OpenStudy (anonymous):

Using a reference triangle, find the exact value of the expression cos(tan^-1x^2)

OpenStudy (anonymous):

tan = opposite/adjacent = x²/1

OpenStudy (anonymous):

Find the third side of the triangle then use it to find cos Θ

OpenStudy (anonymous):

I got \[\sqrt{x^4?}+1\] ?

OpenStudy (anonymous):

I think i found the answer its 1/ the equation from earlier. Thanks alot!!

OpenStudy (anonymous):

\[\cos \theta=\frac{ 1 }{\sqrt{x^4+1} }\]